OpenStudy (babynini):

Convert polynomial to rectangular coordinates (use variables x and y as needed.) r=8+cos(theta)

OpenStudy (babynini):

@Astrophysics ? :)

OpenStudy (babynini):

@zepdrix guess who!

OpenStudy (babynini):

@jim_thompson5910

OpenStudy (babynini):

so I multiplied both sides by r to ge r^2=8r+rcos(theta)

OpenStudy (babynini):

and then (r^2-rcos(theta)^2=8r^2

OpenStudy (babynini):

so \[(x^2+y^2-x)^2=(8r)^2\]

jimthompson5910 (jim_thompson5910):

I see how you multiply both sides by r, but I don't see how you got to step 2

OpenStudy (babynini):

I just moved rcos(theta) to the left hand side

jimthompson5910 (jim_thompson5910):

but you have 8r in step 1, then you have 8r^2 in step 2

OpenStudy (babynini):

\[r^2=8r+rcos \theta\] \[(r^2-rcos \theta) = 8r\] \[(r^2-rcos \theta)^2=(8r)^2\]

jimthompson5910 (jim_thompson5910):

I see. So in step 2 you meant to say (8r)^2 and not 8r^2

OpenStudy (babynini):

Yeah, sorry about that :)

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (babynini):

\[(x^2+y^2-x)^2=(8r)^2\] em..I'm not sure where to go from here

jimthompson5910 (jim_thompson5910):

well r = sqrt(x^2 + y^2) so r^2 = x^2 +y^2 (8r)^2 = 64r^2 = 64*(x^2 +y^2)

jimthompson5910 (jim_thompson5910):

What's strange is that I'm getting a circle but there are going to be x^4 and y^4 terms

OpenStudy (babynini):

hm..I don't know o.0

OpenStudy (babynini):

Well but (x^2+y^2-x)^2=64(x^2+y^2) would be the final thing yeah? I'm not supposed to graph it, phew

jimthompson5910 (jim_thompson5910):

yes that works. We can't solve for y because of the y^2 and y^4 terms

jimthompson5910 (jim_thompson5910):

so it's best to leave it in that implicit form

OpenStudy (babynini):

Yay! it worked, fabulous :) I had been putting 16 instead of 64 *facepalm. Thank youu

jimthompson5910 (jim_thompson5910):

you're welcome